How many pounds of seed worth $0.60 a pound must be mixed with 300 pounds of seed worth $0.35 a pound to get a mixture worth $0.50 per pound
Answer by oscargut(686)

How many pounds of seed worth $0.60 a pound must be mixed with 300 pounds of seed worth $0.35 a pound to get a mixture worth $0.50 per pound
let name x the answer
total cost is 0.6x+0.35(300)
and total pounds are x+300
then (0.6x+105)/x+300 =0.5
then 0.6x+105=0.5x+150
then 0.1x+105=150
then 0.1x=45
then x=450